# Small Stickers, Big Meanings: A Multilingual Sticker Semantic Understanding Dataset with a Gamified Approach

### v1

@@ Abstract.
Stickers, though small, are a highly condensed form of visual expression, ubiquitous across messaging platforms and embraced by diverse cultures, genders, and age groups. Despite their popularity, sticker retrieval remains an underexplored task due to the significant human effort and subjectivity involved in constructing high-quality sticker query datasets. Although large language models (LLMs) excel at general NLP tasks, they falter when confronted with the nuanced, intangible, and highly specific nature of sticker query generation.
To address this challenge, we propose a threefold solution. First, we introduce Sticktionary, a gamified annotation framework designed to gather diverse, high-quality, and contextually resonant sticker queries. Second, we present StickerQueries, a multilingual sticker query dataset containing 1,115 English and 615 Chinese queries, annotated by over 60 contributors across 60+ hours. Lastly, Through extensive quantitative and qualitative evaluation, we demonstrate that our approach significantly enhances query generation quality, retrieval accuracy, and semantic understanding in the sticker domain. To support future research, we publicly release our multilingual dataset [^1] https://huggingface.co/datasets/metchee/sticker-queries along with two fine-tuned query generation models [^2] https://huggingface.co/metchee/sticker-query-generator-en [^3] https://huggingface.co/metchee/sticker-query-generator-zh.
@@
journalyear: 2025conference: Under Review.; ;
@@ 1. Introduction
    Through extensive experiments, we demonstrate significant improvements in sticker retrieval and query generation performance, underscoring the effectiveness of our dataset and methodology.
@@ 2. Related Work
Table 1. Comparison of sticker query datasets sorted by year. Datasets supporting English queries and publicly available are highlighted in bold. Until now, no sticker intention dataset has been publicly available except our proposed StickerQueries.
Dataset Year Description Pub. Avail? Have Queries? Query Language? Native English Query?
Sticker Dialogues
SRS, PESRS (Gao et al., 2021, 2020) 2020 Sticker dialogues.
@@
STICKERCONV(Association for Computational Linguistics 2024 and Zhang, 2024) 2024 Simulated sticker dialogues.
@@
Sticker Sentiment Analysis
CSMSA (Ge et al., 2022) 2022 Sticker sentiment.
@@
Basic Sticker Collections
@@
TGIF (Li et al., 2016) 2016 GIF captioning.
VSD2M (Yuan et al., 2025) 2025 Animated sticker dataset.
Sticker Queries
StickerTag (Wang et al., 2024) 2024 Sticker-tag. Unknown Unknown Unknown
PerSRV (Chee et al., 2024) 2024 Sticker search query. ZH
StickerQueries 2025 Sticker search query. EN, ZH
With the onset of instant messaging, we have observed significant advancements in sticker dataset retrieval. However, given the complexity and diversity of sticker tasks, existing datasets still represent only the tip of the iceberg. We broadly categorize these datasets into four groups: sticker dialogues, sticker sentiment analysis, basic sticker collections, and sticker queries.
Sticker dialogues typically consist of truncated conversations designed to support sticker retrieval or recommendation (Gao et al., 2021, 2020; Fei et al., 2021; Liang et al., 2024; Shi and Kong, 2024; Association for Computational Linguistics 2024 and Zhang, 2024; Wang et al., 2025), whereas U-Sticker (Chee et al., 2025) uniquely provides continuous user information, temporal context, and domain-specific sticker conversations. Among these, several datasets are publicly available, while others remain private or unavailable.
Sticker sentiment datasets focus on predicting the sentiment conveyed by stickers, either as binary positive/negative labels (Ge et al., 2022) or more fine-grained classifications (Liu et al., 2022). Some, like Sticker820K (Zhao et al., 2023), also include supplementary information such as OCR text and descriptive metadata. Within this category, SER30K is publicly accessible, while others are not.
Basic sticker collections mainly provide raw sticker images or GIFs for general usage or generation tasks. Examples include ChineseB2B (Zhaoolee, 2024), which offers a large collection of sticker images, TGIF (Li et al., 2016) for GIF captioning, and the recently introduced VSD2M (Yuan et al., 2025), an animated sticker dataset intended for sticker generation.
Finally, sticker queries datasets are critical as they capture the underlying intention behind sticker usage through condensed, highly precise keywords or search queries. Unfortunately, datasets like StickerTag (Wang et al., 2024) and PerSRV (Chee et al., 2024)are not publicly available for analysis. Though PerSRV (Chee et al., 2024) provides an open source model, their training is limited purely to Chinese queries, restricting its applicability to English-language research. With contrasting cultural nuances, naive translation between languages fails to adequately capture sticker usage intentions. Our StickerQueries dataset addresses these gaps by providing bilingual (English and Chinese) queries crafted by native speakers, thereby bridging the semantic understanding of stickers across diverse cultural contexts and enabling more effective retrieval and recommendation systems.
2.1. Sticker Semantic Understanding
@@ 2.2. Image Captioning
Recent models like BLIP and BLIP-2 (Li et al., 2022, 2023a), as well as GIT (Wang et al., 2022a) and OFASys (Wang et al., 2022b), have advanced image captioning through large-scale vision-language pretraining and unified generative modeling. VinVL (Zhang et al., 2021) and PaLI (Chen et al., 2023) further improve visual representations and multilingual capabilities. While these models produce accurate and fluent descriptions, they focus mainly on tangible visual semantics. In contrast, sticker search revolves around emotional expression, which traditional captioning models are not designed to capture, making them inadequate for emotion-driven query generation.
@@ 3.2. Dataset Goals
  • Multilingual Coverage: Move beyond direct translations to include culturally and contextually appropriate expressions.
  • High Quality Representative Queries: Capture expressions used by native speakers, including local slang and region-specific terms (e.g., “savage”).
  • Diverse Expressions: Reflect varied phrasing for similar intents (e.g., “haha,” “lol,” “lmao”).
@@ 4.1. Game Design
Labeler In each round, the labeler is shown a sticker and asked imagine themselves searching for the sticker, in which they they are expected to key in search queries. Then, their queries are used to search and a preliminary sticker result is shown. This imaginative state prevents player from merely describing the features of the sticker. Based on this, they may revise their queries accordingly. Players are also instructed to avoid repetitive search queries, and purely visual descriptions.
@@
Ensuring Quality Sticktionary’s alternating roles naturally sustain attention and reduce fatigue, promoting higher quality annotations. The point-based system turns precision into reward which further motivates players to craft high quality annotations.
Ensuring Resonance Intuitively, if one player’s query guides another to the correct sticker, it’s a strong indicator of shared understanding and a resonant query. In addition, real-time feedback loops nudge players toward phrasing that feels universally intuitive, rather than overly niche or idiosyncratic.
Ensuring Diversity Lastly, the involvement of more than one round of annotation can ensure diversity and variety. Likewise, if queries can be effectively retrieved, their quality can be ensured.
These mechanisms function in tandem rather than isolation. The built-in game design’s role-switching, feedback loops, and incentive structures collectively ensure that the collected annotations are not only of high quality, but also diverse and widely resonant.
@@ 4.2. Game Settings
4.2.1. Player Recruitment
We recruit participants by incentivizing top performers with themed prizes featuring the university’s mascot panda, along with a participation reward of 30 RMB for completing an observable annotation session.
4.2.2. Data Selection
To ensure the goal of natural sticker associations, we select messages preceded by at least twenty words of dialogue, filtering out command-driven interactions and exchanges dominated by brief replies. This ensures that the stickers appear in expressive, emotionally grounded contexts. We apply these filters to the U-Sticker dataset (Chee et al., 2025), followed by a manual review. The final collection consists of 1,136 non-duplicated conversations, each curated for contextual richness and intuitive alignment.
@@ 4.3. Game Statistics
Table 2. Summary of player participation and task distribution across English and Chinese versions of the game after outlier removal. The dataset reflects contributions from over 62 players and approximately 60 hours of gameplay.
@@ 5.1. Statistics
Table 3. StickerQuery comprises over 1,700 multilingual sticker-query pairs, each annotated by at least two annotators, offering a rich, diverse set of natural expressions.
@@
Table 3 presents overview of the StickerQuery dataset statistics, highlighting its scale, linguistic diversity, and annotation coverage. For the following steps, we use Jieba for Chinese tokenization (Junyi, 2012).
Figure 4. Word cloud of the English StickerQueries, feelings dominate the visual.
Figure 5. Frequency distribution of the top 10 most common English query words in the StickerQueries dataset.
We find that emotional expressions dominate sticker queries: 9 of the top 10 entries in both English and Chinese pertain to feelings rather than actions or visuals (Figures 5 and 7). This aligns with the idea that users are often driven by internal emotional states when searching, rather than external depictions.
Interestingly, although survey participants emphasized action-oriented queries, annotation patterns reveal a stronger inclination toward emotional resonance—affirming the sticker’s role as an affective medium. While English and Chinese users show largely parallel emotional vocabularies, terms like lol, chill, and zen emerge as uniquely English, suggesting subtle cultural inflections in how emotion is expressed across languages.
Figure 6. Word cloud of the Chinese StickerQueries, feelings dominate the visual.
Figure 7. Top 10 most common Chinese query words in the StickerQueries dataset, showing the expressions trends.
@@ 5.2. Interannotator Similarity
Table 4. Multilingual sticker query similarity evaluation. BLEU and ROUGE are low due to short queries, while BERT-based and cosine similarity show higher semantic alignment.
Metric En Zh
BLEU (Papineni et al., 2002): 0.0756 0.0132
ROUGE-2: 0.1524 0.0034
ROUGE-L: 0.2796 0.0195
CosineSim: 0.7131 0.7728
BERTScore (P) (Zhang et al., 2020) 0.6710 0.7061
BERTScore (R) 0.6666 0.6983
BERTScore (F1) 0.6628 0.6986
We use the pretrained BERT-base models released by Google for English (Devlin et al., 2019) and Chinese to compute both cosine similarity and BERTScore (Zhang et al., 2020) between query annotations. ROUGE-1, ROUGE-2, and ROUGE-L (Lin, 2004) are also reported for comparison with traditional string-based metrics.
5.2.1. English Annotations
English annotations show higher lexical overlap, with BLEU (0.0756) and ROUGE scores (e.g., ROUGE-1: 0.3019), indicating strong n-gram similarity. Semantic metrics also confirm good meaning alignment (cosine: 0.7131, BERT F1: 0.6628).
5.2.2. Chinese Annotations
Chinese annotations score lower on lexical metrics (BLEU: 0.0132, ROUGE-1: 0.0195), likely due to greater paraphrasing. However, they achieve higher semantic similarity (cosine: 0.7728, BERT F1: 0.6986), indicating strong meaning preservation despite fewer surface overlaps.
Following prior work (Zhang et al., 2020), a BERTScore F1 above 0.65 indicates moderate semantic agreement in open-ended generation tasks. Our dataset achieves 0.66–0.70, validating the reliability of annotations while retaining expressive diversity.
@@ 6.1. Query Generation
We split the data into 20% test and 80% training sets and fine-tuned LLaVA-1.5B-7B, a moderately sized vision-language model, on our StickerQueries dataset.
Table 5. Performance of LLaVA-1.5-7B on StickerQueries, showing zero-shot vs fine-tuned results. Significant improvements demonstrate effective adaptation to both English and Chinese query generation.
Model BLEU ROUGE-2 ROUGE-L CosSim
English Evaluation
Zero-Shot 11.08 2.82 19.22 0.778
Fine-Tuned (SFT) 15.62 8.26 27.05 0.811
Improvement $\uparrow$40.9% 192.5% 40.8% 4.2%
Chinese Evaluation
Zero-Shot 0.5138 0.0296 2.2584 0.6349
Fine-Tuned (SFT) 6.3432 2.970 24.26 0.7976
Improvement $\uparrow$1134.6% 9934.5% 974.2% 25.6%
Fine-tuning our StickerLabel dataset leads to clear improvements across all evaluation metrics for both English and Chinese, in particular significant gains in ROUGE scores and cosine similarity. This demonstrates the dataset’s effectiveness in enhancing the model’s ability to generate accurate, contextually relevant sticker queries.
@@ 6.2. Sticker Retrieval
R@1/50 $\Delta\%$ R@1/50 $\Delta\%$
Context 0.018/0.297 900/129 0.008/0.407 413/12
BLIP2 0.032/0.248 462/174 0.005/0.225 720/102
Vanilla VLM 0.153/0.644 18/6 0.016/0.423 156/8
SFT VLM 0.180/0.680** 0.041/0.455**
Since we used a 20% test split, we compare three sticker semantic baseline methods for the sticker retrieval task to generate queries.
  • Context: Take the previous utterance in the context as the query.
@@
  • Vanilla VLM(Liu et al., 2023): Zero-shot query generation with LLaVA-1.5B-7B.
  • SFT VLM (Ours): Finetuned LLaVA-1.5B-7B model on StickerQueries dataset (Liu et al., 2023).
@@
7. Discussion and Future Work
We understand that our dataset is limited and expanding it would likely enhance performance and robustness. Additionally, incorporating more multilingual data and increasing the number of survey respondents could help us to understand sticker usage behavior.

### v2

@@ Abstract.
Stickers, though small, are a highly condensed form of visual expression, ubiquitous across messaging platforms and embraced by diverse cultures, genders, and age groups. Despite their popularity, sticker retrieval remains an underexplored task due to the significant human effort and subjectivity involved in constructing high-quality query-sticker datasets. Although large language models (LLMs) excel at general NLP tasks, they falter when confronted with the nuanced, intangible, and highly specific nature of sticker query generation.
To address the challenge of collecting diverse and contextually appropriate sticker search queries, we introduce Sticktionary, a gamified annotation framework designed to elicit high-quality, semantically rich queries from contributors. Using this framework, we construct StickerQueries, a multilingual dataset comprising 1,115 English and 615 Chinese sticker search queries, annotated by over 60 contributors across more than 60 hours of annotation. We demonstrate the utility of StickerQueries in several downstream tasks, including query generation and sticker retrieval. Through comprehensive quantitative and qualitative evaluations, we demonstrate that StickerQueries significantly improves the quality of query generation, retrieval accuracy, and semantic understanding within the sticker domain. To support future research, we publicly release the multilingual StickerQueries dataset and two fine-tuned query generation models [^1] https://huggingface.co/datasets/metchee/sticker-queries. Additional experiments and supplementary case studies can be found here  [^2] https://arxiv.org/abs/2506.01668.
@@
journalyear: 2025conference: the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13-18, 2025; Padua, Italyjournalyear: 2025copyright: ccconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3758310isbn: 979-8-4007-2035-2/2025/10ccs: Information systems Information retrievalccs: Information systems Dataset
@@ 1. Introduction
    Through extensive experiments, we demonstrate significant improvements in sticker retrieval and query generation, validating our dataset and methodology.
@@ 2. Related Work
Table 1. Comparison of sticker datasets sorted by characteristics and year. Currently, no sticker query dataset has been publicly available except our proposed StickerQueries. We use NA as a shorthand for Not Available data values.
Category Dataset Year Description Publicly Available? Query Language? Native EN?
Dialogues SRS, PESRS (Gao et al., 2021, 2020) 2020 Sticker dialogues.
@@
STICKERCONV(Zhang et al., 2024) 2024 Simulated sticker dialogues.
@@
Sentiment CSMSA (Ge et al., 2022) 2022 Sticker sentiment.
@@
Captions TGIF (Li et al., 2016) 2016 GIF captioning.
StickerTag (Wang et al., 2024) 2024 Sticker-tag. NA NA
Multimedia VSD2M (Yuan et al., 2025) 2025 Animated sticker dataset.
@@
Retrieval PerSRV (Chee et al., 2024) 2024 Sticker-query pairs. ZH
StickerQueries (Ours) 2025 Sticker-query pairs. EN, ZH
2.1. Sticker Semantic Datasets
With the onset of instant messaging, we have observed significant advancements in sticker dataset retrieval. However, given the complexity and diversity of sticker tasks, existing datasets still represent only the tip of the iceberg. We broadly categorize these datasets into five groups: dialogues, sentiment, captions, multimedia and sticker queries.
Dialogue datasets typically consist of truncated conversations containing stickers (Gao et al., 2021, 2020; Fei et al., 2021; Liang et al., 2024; Shi and Kong, 2024; Association for Computational Linguistics 2024 and Zhang, 2024; Wang et al., 2025). In contrast, U-Sticker (Chee et al., 2025) uniquely provides continuous user information, temporal context, and domain-specific sticker conversations. While some of these datasets are publicly available, others remain private or inaccessible. Sentiment datasets focus on predicting the sentiment conveyed by stickers, either as binary positive/negative labels (Ge et al., 2022) or through more fine-grained classifications (Liu et al., 2022). Certain datasets, such as Sticker820K (Zhao et al., 2023), also include supplementary information like OCR text and descriptive metadata. Within this category, SER30K (Liu et al., 2022) is publicly available, whereas others are not. Caption datasets contain stickers with textual elements. For instance, TGIF (Li et al., 2016) provides captions for GIFs. Unfortunately, StickerTag (Wang et al., 2024) is unavailable, preventing further analysis. Multimedia datasets include raw sticker images for general use or generation tasks. ChineseB2B (Zhaoolee, 2024) offers a large collection of sticker images and VSD2M (Yuan et al., 2025) contains an extensive repository of sticker animations.
Finally, retrieval datasets pair user search queries with corresponding sticker images. These datasets are critical for capturing the underlying intent behind sticker usage through concise, high-precision keywords. Unfortunately, the only other sticker–query pair dataset, PerSRV (Chee et al., 2024), is not publicly available for analysis. Although PerSRV provides an open-source model, its training is limited to Chinese queries, restricting its applicability to English-language research. Furthermore, cultural differences mean that naive translation between languages often fails to convey the true intent of sticker usage. Our StickerQueries dataset addresses these limitations by offering bilingual (English and Chinese) queries crafted by native speakers, thereby bridging semantic understanding across cultural contexts and enabling more effective retrieval and recommendation systems.
@@ 2.2. Image Captioning
Recent models like BLIP and BLIP-2 (Li et al., 2022, 2023a), as well as GIT (Wang et al., 2022a) and OFASys (Wang et al., 2022b), have advanced image captioning through large-scale vision-language pretraining and unified generative modeling. VinVL (Zhang et al., 2021) and PaLI (Chen et al., 2023) further improve visual representations and multilingual capabilities. While these models produce accurate and fluent descriptions, they focus mainly on tangible visual semantics. In contrast, sticker search queries mainly revolves around emotional expression, which traditional captioning models are not designed to capture, making them inadequate for emotion-driven query generation.
@@ 3.2. Dataset Goals
  • Multilingual Coverage: Avoid naive translations by incorporating various languages.
  • High Quality Representative Queries: Capture expressions used by native speakers, including local slang and region-specific terms.
  • Diverse Expressions: Reflect varied phrasing for similar intents (e.g., “haha,” “laugh out loud”).
@@ 4.1. Game Design
Labeler In each round, the labeler is shown a sticker and asked imagine themselves searching for the sticker, in which they are expected to key in search queries. Then, their queries are used to search and a preliminary sticker result is shown. This imaginative state prevents player from merely describing the features of the sticker. Based on this, they may revise their queries accordingly. Players are also instructed to avoid repetitive search queries, and purely visual descriptions.
@@ 4.2. Game Settings
Player Recruitment. We recruit participants by incentivizing top performers with themed prizes featuring the university’s mascot panda, along with a participation reward of 30 RMB for completing an observable annotation session.
Data Selection. To ensure the goal of natural sticker associations, we select messages preceded by at least twenty words of dialogue, filtering out command-driven interactions and exchanges dominated by brief replies. This ensures that the stickers appear in expressive, emotionally grounded contexts. We apply these filters to the U-Sticker dataset (Chee et al., 2025), followed by a manual review. The final collection consists of 1,136 non-duplicated conversations, each curated for contextual richness and intuitive alignment.
4.3. Goal-orientated
Ensuring High Quality Queries. Sticktionary’s alternating roles sustain attention and reduce fatigue, boosting annotation quality. Points reward precise queries that retrieve the correct sticker, naturally validating their effectiveness. This retrieval-based gameplay also provides a natural validation signal, as successful retrievals indicate effective and contextually appropriate queries.
Ensuring Resonance. Multiple successful retrieval from also indicates shared understanding and a resonant query. In addition, real-time feedback loops nudge players toward phrasing that feels universally intuitive, rather than overly niche or idiosyncratic.
Ensuring Diversity. Lastly, the involvement of more than one round of annotation can ensure diversity and variety. Each participant brings their own linguistic and cultural lens, contributing to a rich set of expressions for the same sticker. Furthermore, the reliance on retrievability as a success criterion naturally filters out repetitive or low-impact queries, ensuring that only distinct and effective expressions are retained.
Ensuring Multilingual Coverage, Representative Queries and Natural Associations. Players participate exclusively in their native language, ensuring authentic and culturally grounded expressions. Our data filtering process ensures that the stickers appear in expressive, emotionally grounded contexts.
These mechanisms work together—role-switching, feedback, and incentives—to ensure annotations are high-quality, diverse, and broadly resonant.
@@ 4.4. Game Statistics
Table 2. Summary of player participation and task distribution across English and Chinese versions of the game after outlier removal. The dataset reflects contributions from over 60 players and approximately 60 hours of gameplay.
@@ 5.1. Statistics
Table 3. StickerQueries comprises over 1,700 multilingual sticker-query pairs, each annotated by at least two annotators, offering a rich, diverse set of natural expressions.
@@
Table 3 presents overview of the StickerQueries dataset statistics, highlighting its scale, linguistic diversity, and annotation coverage. For the following steps, we use Jieba for Chinese tokenization (Junyi, 2012).
Figure 4. Word cloud of the English (top) and Chinese (bottom) StickerQueries, feelings dominate the visual.
Figure 5. Frequency distribution of the top 10 most common English (top) and Chinese (bottom) query words in the StickerQueries dataset.
From Figures 5 and 4, we observe that emotional expressions dominate sticker queries: 9 of the top 10 terms in both English and Chinese relate to feelings rather than actions or visuals. This supports the view that users are often guided by internal emotional states when searching, rather than external depictions. Although survey participants emphasized action-oriented queries (Fig  2, annotation patterns preliminarily reveals a stronger tendency toward emotional resonance—affirming stickers’ role as an affective communication medium.
Additionally, while English and Chinese users share many emotional sentiments, subtle cross-linguistic and cultural differences emerge. English-exclusive terms include lol, chill, and zen, whereas culturally rooted Chinese expressions like eating melon (watching gossip), do rice (eating a meal), and awesome awesome one (awesome) appear. These highlight the limitations of naive translation and emphasize the cultural richness captured in the StickerQueries dataset
@@ 5.2. Interannotator Similarity
Table 4. Multilingual sticker query similarity evaluation. BLEU and ROUGE (R) are low due to short queries, while BERT-based (B) and cosine similarity (CS) show higher semantic alignment.
BLEU R-L CS B (P) B (R) B (F1)
EN 0.0756 0.2796 0.7131 0.6710 0.6666 0.6628
ZH 0.0132 0.0195 0.7728 0.7061 0.6983 0.6986
We use the pretrained BERT-base models (Devlin et al., 2019) released by Google for English and Chinese to compute both cosine similarity and BERTScore (Zhang et al., 2020) between query annotations. ROUGE-L (Lin, 2004) are also reported for comparison with traditional string-based metrics.
English annotations show higher lexical overlap, with BLEU (0.0756) and ROUGE scores (e.g., ROUGE-L: 0.2796), indicating strong n-gram similarity. Semantic metrics also confirm good meaning alignment (cosine: 0.7131, BERT F1: 0.6628). Chinese annotations score lower on lexical metrics (BLEU: 0.0132, ROUGE-L: 0.0195), likely due to greater paraphrasing. However, they achieve higher semantic similarity (cosine: 0.7728, BERT F1: 0.6986), indicating strong meaning preservation despite fewer surface overlaps. Following prior work (Zhang et al., 2020), a BERTScore F1 above 0.65 indicates moderate semantic agreement in open-ended generation tasks. Our dataset achieves 0.66–0.70, validating the reliability of annotations while retaining expressive diversity.
@@ 6.1. Query Generation
We split the data into 80% training and 20% validation and test sets, then fine-tuned LLaVA-1.5B-7B on our StickerQueries dataset. Fine-tuning on StickerQueries dataset leads to clear improvements across all evaluation metrics for both English and Chinese, as shown in Table  5. This demonstrates the dataset’s effectiveness in enhancing the model’s ability to generate accurate, contextually relevant sticker queries.
Table 5. Performance of LLaVA-1.5-7B on StickerQueries, showing zero-shot vs fine-tuned results. Significant improvements demonstrate effective adaptation to both English and Chinese query generation. R short for ROUGE.
Model BLEU R-2 R-L CS
EN Vanilla VLM 11.08 2.82 19.22 0.78
SFT VLM 15.62 8.26 27.05 0.81
ZH Vanilla VLM 0.51 0.03 2.26 0.63
SFT VLM 6.34 2.97 24.26 0.80
@@ 6.2. Sticker Retrieval
R@1 R@50 R@1 R@50
Context 0.018 0.297 0.008 0.407
BLIP2 0.032 0.248 0.005 0.225
Vanilla VLM 0.153 0.644 0.016 0.423
SFT VLM 0.180 0.680 0.041 0.455
On the same 20% test split, we compare three sticker semantic baseline methods for the sticker retrieval task to generate queries.
  • Context: Take previous utterance as query.
@@
  • Vanilla VLM (Liu et al., 2023): Zero-shot LLaVA-1.5B-7B.
  • SFT VLM (Ours): Finetuned LLaVA-1.5B-7B model.
@@
7. Limitations and Future Work
We acknowledge our dataset’s limitations, including a primarily young adult annotation pool and a relatively small size. Expanding it with more survey respondents, diverse participants, additional annotations, and multilingual data would likely improve performance, robustness, and understanding of sticker usage behavior.